OpenStudy (anonymous):

11. Solve the inequality. Show your work. |r + 3| ≥ 7

OpenStudy (anonymous):

i need help

OpenStudy (rulnick):

The statement |r+3| >= 7 is equivalent to "r+3 >=7 or r+3 <= -7"

OpenStudy (rulnick):

So if you solve these two inequalities separately and join their answers with "or" you'll have it right.